OpenStudy (anonymous):

I need help with finding the intercepts of these equations y=X^2(sqrt( 25-X^2) and y=3(2- sqrt (x))/over x

myininaya (myininaya):

you want where the equations intersect? or do you really want the intercepts of each?

OpenStudy (anonymous):

i want like x intercepts and y and how you re-arrange the equation

myininaya (myininaya):

to find x-intercept(s), set y=0 and solve for x. to find y-intercept(s), set x=0 and solve for y

myininaya (myininaya):

so for example: we have \[y=x^2\sqrt{25-x^2}\]. to find x-intercepts set y=0 and solve for x. \[0=x^2\sqrt{25-x^2}\]

myininaya (myininaya):

\[x^2=0, \sqrt{25-x^2}=0\]

myininaya (myininaya):

\[x=0, x=\pm 5\]

myininaya (myininaya):

you have 3 x-intercetps (0,0) (5,0) (-5,0)

OpenStudy (anonymous):

oooh can you pls do the next one the square roots and division sign confuse m when settin them = to 0

myininaya (myininaya):

now to find the y-intercept set x=0 and solve for y \[y=0^2\sqrt{25-0^2}=0\sqrt{25}=0\] so we have y-intercept is (0,0) but could have prevented doing this step since we already found the y-intercept above

myininaya (myininaya):

for the other equation?

myininaya (myininaya):

oops one second

myininaya (myininaya):

to find y-intercepts set x=0but what happens when you plug 0 in where x is?

OpenStudy (anonymous):

you mean y = 0 to find x intercepts?

OpenStudy (anonymous):

the square root disapear?

myininaya (myininaya):

don't you have x in the denominator?

OpenStudy (anonymous):

oh ya

myininaya (myininaya):

so the function is undefined when x=0 so no y-intercepts

myininaya (myininaya):

now to find x-intercepts, set y=0 and solve for x

OpenStudy (anonymous):

i c, can you showme steps when setting 6-3(squrt x)=0

OpenStudy (anonymous):

like how you move the 3 and so on

myininaya (myininaya):

a fraction is only zero when the stop is zero so we are looking at solving \[0=3(2-\sqrt{x})\]

myininaya (myininaya):

divide both sides by 3

myininaya (myininaya):

then add square root of x on both sides

myininaya (myininaya):

and then square both sides

OpenStudy (anonymous):

oooh thanks can you show sqrt part so i can c visuallysrory

OpenStudy (anonymous):

then show x= and x= so i can check

myininaya (myininaya):

show me what you think i just asked you to do and i will be right back i have to get to ready to go to sleep

OpenStudy (anonymous):

ok so 3=0 and (2-squrtX=0?

OpenStudy (anonymous):

ok so 3=0 and (2-squrtX=0?

myininaya (myininaya):

well 3 is not equal to 0 so we have \[2-\sqrt{x}=0\] add \[\sqrt{x}\] on both sides

myininaya (myininaya):

then square both sides

OpenStudy (anonymous):

oh could u also have subtracted 2?